OpenStudy (anonymous):

Solve this DE:

OpenStudy (anonymous):

\[\sqrt{1-y^2}dx-\sqrt{1-x^2}dy=0\]

OpenStudy (anonymous):

\[\sqrt{1-y^2}dx=\sqrt{1-x^2}dy\]

OpenStudy (anonymous):

\[\frac{dx}{\sqrt{1-x^2}}=\frac{dy}{\sqrt{1-y^2}}\]

OpenStudy (ash2326):

We can write this as \[\frac{dy}{\sqrt{1-y^2}}=\frac{dx}{\sqrt{1-x^2}}\] Now integrate both sides \[\int\frac{dy}{\sqrt{1-y^2}}=\int \frac{dx}{\sqrt{1-x^2}}\] This is a standard integral, can you do this ?

OpenStudy (anonymous):

yeah it's arcsin however it asks what specific solution at y(square root of 3 over 2) arcsin(y)=arcsin(x)+c should i multiply both sides by sin to cancel and get in y or should i just keep this like so

OpenStudy (anonymous):

y=x+sin(c)

OpenStudy (anonymous):

i guess i should keep since i'd have to take the arcsin to get back to c

OpenStudy (ash2326):

Is the value of x given?

OpenStudy (anonymous):

yeah y(0)= sqrt 3 / 2

OpenStudy (ash2326):

put x=0 \[\sin^{-1} \frac{\sqrt 3}{2}=\sin^{-1}(0)+c\] \[\frac{\pi}{3}=0+c\] Now our DE is \[\sin^{-1} y=\sin^{-1}x+\frac{\pi}{3}\]